# Cross-Lingual Transfer of Debiasing and Detoxification in Multilingual LLMs: An Extensive Investigation

### v2

@@ Abstract
Recent generative large language models (LLMs) show remarkable performance in non-English languages, but when prompted in those languages they tend to express higher harmful social biases and toxicity levels. Prior work has shown that finetuning on specialized datasets can mitigate this behavior, and doing so in English can transfer to other languages. In this work, we investigate the impact of different finetuning methods on the model’s bias and toxicity, but also on its ability to produce fluent and diverse text. Our results show that finetuning on curated non-harmful text is more effective for mitigating bias, and finetuning on direct preference optimization (DPO) datasets is more effective for mitigating toxicity. The mitigation caused by applying these methods in English also transfers to non-English languages. We find evidence that the extent to which transfer takes place can be predicted by the amount of data in a given language present in the model’s pretraining data. However, this transfer of bias and toxicity mitigation often comes at the expense of decreased language generation ability in non-English languages, highlighting the importance of developing language-specific bias and toxicity mitigation methods.
@@ 1 Introduction
Most generative large language models (LLMs) are primarily trained on English data (Jiang et al., 2023; Hayase et al., 2024), though they are used by speakers of many different languages (Zheng et al., 2024). While these models show remarkable performance in non-English languages Shi et al. (2023), they also express harmful social biases and high toxicity levels when prompted in those languages (Jain et al., 2024; Üstün et al., 2024; Wang et al., 2024b), as well as in English (Nangia et al., 2020; Nadeem et al., 2021; Gehman et al., 2020; Zheng et al., 2024; Lin et al., 2023).
@@
Figure 1: Disclaimer: Potentially sensitive content. An example of a toxic generation by Llama 3.1 Instruct for a German prompt from RTP-LX. After finetuning on DetoxDPO the generation is no longer toxic, but also no longer in German.
@@
To mitigate this behavior, LLMs are finetuned to provide safe answers to harmful user inputs (Üstün et al., 2024; Allam, 2024). Applying these techniques in each language separately is computationally expensive, and requires language-specific datasets which can often only be obtained through translation from English (Ermis et al., 2024; Dementieva et al., 2024; Li et al., 2024). Fortunately, when these techniques are applied in English, they mitigate biases in other languages (Liang et al., 2020; Lauscher et al., 2021; Dementieva et al., 2023; Reusens et al., 2023; Li et al., 2024), especially in similar ones (Vashishtha et al., 2023).
@@
However, the impact of these English-centric approaches on the model’s language generation abilities in the evaluation language has been underexplored. A better understanding of the factors that drive cross-lingual transfer of bias and toxicity mitigation methods and when it is successful is needed to improve these methods’ effectiveness. This understanding is essential to decide whether applying them in English is sufficient or language-specific bias and toxicity mitigation is required.
@@
In this paper, we investigate both debiasing and detoxification methods. Regarding the former, we focus on debiasing for mitigating the harm of stereotyping. Stereotypes are overgeneralized beliefs about an individual’s personal characteristics based on the demographic group to which they belong (Greenwald and Banaji, 1995; Dev et al., 2022), e.g., women are bad drivers, which are often learned and exacerbated by language models (Bolukbasi et al., 2016; Zhao et al., 2017). Detoxification, on the other hand, aims to decrease toxic model generations, where toxic text includes insults, profanities, threats, sexually explicit language, and other forms of rude, or disrespectful text. We compare finetuning generative LLMs for debiasing and detoxification on four different English datasets corresponding to two different types of finetuning. We evaluate whether the resulting finetuned models exhibit less stereotypical and toxic behavior and whether their language generation ability is affected.
Concretely, we aim to answer the following research questions:
@@
    Which finetuning methods and existing datasets for mitigating bias and toxicity in English are effective to mitigate biased and toxic behavior also in other languages?
@@
Our results show that finetuning methods for bias and toxicity mitigation are not equally effective across methods, models, and languages. Whenever debiasing and detoxification succesfully transfer from English, this comes paired with a decrease in language generation ability in other languages, as illustrated in Figure 1. Successful transfer for a specific bias type depends on the types of bias covered by the finetuning dataset, and transfer to a specific language can to some extent be predicted by the amount of pretraining data in that language. Our work highlights the importance of carrying out debiasing and detoxification in the evaluation language, which is why we hope to encourage future development of multilingual or non-English bias and toxicity mitigation datasets. [^1] Our code is available at https://github.com/Veranep/crosslingualdetoxdebias.
@@ 2 Related Work
Previous work on multilingual toxicity found that models trained on synthetic English preference-tuning data are less toxic than those trained on human English preference-tuning data when evaluated in English, but the opposite is true when evaluated in non-English languages (Jain et al., 2024). However, this preference-tuning data was not specifically aimed at addressing model toxicity. Dementieva et al. (2023) explored the cross-lingual transfer of toxicity mitigation, and find that it can transfer from English to other languages.
Recent work by Li et al. (2024) showed that DPO for toxicity mitigation transfers from English to other languages by decreasing activations for a language-agnostic region in the model responsible for toxic generations. Similar to our findings, they found that this transfer comes at the expense of the diversity and fluency of the model’s generations. Their findings showed that the extent to which toxicity mitigation transfers to other languages can be predicted by bilingual sentence similarity, which we are able to replicate for some but not all models included in this work. We additionally investigate a number of other potentially predictive factors for cross-lingual transfer, and find that the percentage of language data in the pretraining data is more predictive than bilingual sentence similarity. Compared to the work by Li et al. (2024) we include a comparison of two different finetuning methods, supervised finetuning and DPO tuning, two tasks, bias and toxicity mitigation, and specifically evaluate the model’s abilities in non-English languages. For non-English languages we observe a considerable decrease in generations that match the language of the prompt and a substantial reduction in quality of generations as a result of English finetuning. Therefore, unlike prior work we suggest to heed caution when relying on cross-lingual transfer of toxicity and bias mitigation.
@@
LLMs forget previously learned information when being trained on new information, a phenomenon known as catastrophic forgetting (McCloskey and Cohen, 1989; Kirkpatrick et al., 2017). This also affects model debiasing and detoxification. Meade et al. (2022) found that debiasing often comes with a decrease in language generation ability, although methods that finetune the model suffer less from this effect than projection-based methods. Welbl et al. (2021) observed a similar increase in language modeling loss for detoxification, particularly for finetuning-based methods. In addition, Fatemi et al. (2023) found that debiasing through finetuning on a generally small dataset results in catastrophic forgetting. To exacerbate the situation further, Liu et al. (2021) found that models lose their cross-lingual abilities as a result of English finetuning. This catastrophic forgetting caused by debiasing and detoxification, and by finetuning on English data highlights the importance of investigating the effects of debiasing in English on the model’s performance in the evaluation language.
Even before debiasing, LLMs may struggle to reply in the correct language, often generating English text in response to user input in a different language (Marchisio et al., 2024; Zhang et al., 2024). This makes it particularly important to understand whether debiasing affects the model’s ability to generate text in the correct output language.
@@ 3.1 Debiasing and Detoxification Datasets
Overall, we use four different datasets, two for each task. All datasets are in English, since our goal is to investigate cross-lingual transfer from English to other languages.
Bias
The SFT dataset we use for bias mitigation is Perturbation Augmentation NLP DAtaset (PANDA) (Qian et al., 2022), a human-annotated dataset consisting of 98K sentence pairs of original and demographically perturbed sentences. The perturbed sentences have been modified along the gender, race/ethnicity, and age axes, perturbing a sentence such as “women like shopping” to “men like shopping”. Fine-tuning on these perturbed sentences has been shown to be an effective debiasing technique (Prakash and Lee, 2023; Ranaldi et al., 2024). For DPO finetuning we use BiasDPO (Allam, 2024), a manually designed DPO dataset for mitigating gender, race, and religion biases. The dataset consists of 1,1K prompts with biased and unbiased completions.
Toxicity
The SFT dataset we use for toxicity mitigation is the English Jigsaw Civil Comments dataset. [^2] https://www.kaggle.com/c/jigsaw-unintended-bias-in-toxicity-classification These comments from the Civil Comments platform are human-annotated for toxicity, and we follow Xu et al. (2021) in selecting those with less than 10% toxic annotations as our non-harmful dataset. We select a subset of 98K sentences, matching the size of the PANDA dataset. Lee et al. (2024) created a DPO dataset specific for toxicity, which has been shown to also mitigate toxicity in languages other than English (Li et al., 2024). The dataset consists of 25K prompts from Wikitext-2 (Merity et al., 2017) with synthetic toxic and non-toxic completions. This is the DPO dataset we use, which we will refer to as DetoxDPO. See Table 2 in Appendix A for an example from each dataset. Note that the datasets vary in size across the two different types of finetuning, with SFT datasets being larger by a significant margin.
@@ 3.2 Models and Finetuning Procedure
We finetune the models on the datasets described in Section 3.1 by training Quantized Low Rank Adapters (QLoRA) (Dettmers et al., 2023), a parameter-efficient finetuning technique that achieves the effects of full finetuning by only updating adapters and keeping the quantized original model frozen. The exact hyperparameters and compute budget used are provided in Appendix C.
@@ 3.3 Evaluation
We evaluate the models before and after finetuning regarding their levels of bias, the toxicity of their generations, and various aspects of their language generation ability. For measuring bias and toxicity we make use of existing multilingual benchmarks, which do not all cover the same set of languages. For measuring language generation ability we include all $20$ languages that occur in the union of the bias and toxicity benchmarks. [^5] Arabic, Catalan, Chinese, Czech, Dutch, English, French, German, Hindi, Indonesian, Italian, Japanese, Korean, Maltese, Polish, Portuguese, Russian, Spanish, Swedish, Turkish.
@@
Language generation ability
As we are interested in the negative impact that finetuning in English may have on LLMs’ abilities in other languages, we record three different language generation ability metrics in each evaluation language: language consistency, fluency, and diversity.
Following Li et al. (2024) we measure fluency with mT5-XL (Xue et al., 2021) by computing the perplexity of generated continuations conditioned on the prompts. We report the inverse perplexity as fluency, since a lower perplexity corresponds to more fluent generations. [^8] In contrast to Li et al. (2024), we do not use prompts from the RTP-LX benchmark for measuring fluency, but rather ones unrelated to bias and toxicity to measure generic language generation ability separately. Like Li et al. (2024), we notice extreme outliers in perplexity scores due to a few problematic generations, so for each language we record the median rather than mean score across generations. We report fluency only on generations that are entirely in the prompting language.
To evaluate whether models generate continuations beyond mere repetitions of the input prompt, we additionally measure the diversity of the model’s generations, defined as the percentage of unigrams that did not occur in the input prompt. Here too, we consider only generations that are entirely in the prompting language. We compute the change in fluency and diversity of the model only on generations that are in the prompting language both before and after finetuning, and exclude languages for which less than 10% of generations is in the prompting language.
@@ 4 Results
We first evaluate all models in Section 4.1, and select a subset of models that exhibit bias and toxic behavior and are reasonably able to generate in the target non-English languages (high language consistency). This subset of models is then finetuned in English with the goal of debiasing and detoxification. Next, in Section 4.2 we discuss the cross-lingual effects of this finetuning on the model’s bias and toxicity levels, and on their language generation abilities in the non-English evaluation languages.
We report the results of our initial evaluation of instruction-tuned models in Table 1 and the results for the base models in Table 6 in Appendix H. Even though we focus our evaluation on the non-English languages included in the benchmarks, we separately report the results for English in Appendix J.
@@ 4.2 Evaluation of Finetuned Models
We now move on to investigate the changes in bias, toxicity and language generation ability as a result of finetuning. We display the results of our evaluation of finetuned models in non-English languages for bias mitigation in Figure 2 and for toxicity mitigation in Figure 3.999We display the results of evaluating base models in non-English languages in Figure 10 and Figure 11 in Appendix H for bias and toxicity mitigation respectively. Again, for completeness we include the same evaluation for English in Figure 12 and Figure 13 for instruction models, and in Figure 14 and Figure 15 in Appendix J for base models.
@@
(c) Language generation abilities
Figure 2: Effects of bias mitigation on bias scores, and language generation scores. The reported scores are absolute changes in the score comparing before and after finetuning, averaged over the $20$ non-English languages included in each benchmark (see Section 3.3). Errorbars indicate 95% confidence intervals.
@@
(b) Language generation abilities
@@
Supervised finetuning on Panda decreases bias on all three bias benchmarks in nearly all cases (Figures 2(a) and 2(b)), and is therefore more effective for bias mitigation than BiasDPO. However, it is also accompanied by with a sharp decrease in ability to generate in the correct language and a decrease in diversity of generations (Figure 2(c)), particularly for models that generated diverse texts before debiasing. Surprisingly, finetuning on the Jigsaw dataset for toxicity mitigation leads to an undesirable increase in toxicity (Figure 3(a)). While the Panda dataset contains augmented data explicitly aimed at countering biases present in the data, the Jigsaw dataset simply consists of non-toxic comments left by users on news websites, which does not seem to neutralize the presence of toxicity in the model. Similar to the Panda dataset, Jigsaw also results in a decrease in ability to generate in the correct language and a decrease in diversity of generations, though the generations in the correct language increase in fluency (Figure 3(b)).
@@
Despite the relatively small dataset size, DPO preference tuning on BiasDPO successfully mitigates bias on all benchmarks for the Aya models (Figures 2(a) and 2(b)). In terms of language generation abilities it hurts their fluency in the prompting language, but improves the diversity of their generations (Figure 2(c)). DPO preference tuning is more effective for English than for other languages, though it is again only for the Aya models that it approaches the effectiveness of supervised finetuning. For toxicity mitigation DPO training on DetoxDPO is more effective than SFT, given that it reduces toxicity in English and non-English languages in all models (Figure 3(a)). Surprisingly, we find that also here the diversity of the model’s generations in the evaluation language increases, in contrast to prior findings that show DPO training reduces diversity (Li et al., 2024; Wang et al., 2024a). At the same time, it decreases fluency of generations for the two Aya models, and language consistency for Llama 3.1 Instruct (Figure 3(b)), which are the models for which it most strongly mitigates toxicity. In Appendix E we visualize how bias, toxicity, and language generation ability scores develop during training, which suggest that effects similar to DetoxDPO’s could have been achieved by BiasDPO if it had been larger.

### v3

@@ Abstract
Recent generative large language models (LLMs) show remarkable performance in non-English languages, but when prompted in those languages they tend to express higher harmful social biases and toxicity levels. Prior work has shown that finetuning on specialized datasets can mitigate this behavior, and doing so in English can transfer to other languages. In this work, we investigate the impact of different finetuning methods on the model’s bias and toxicity, but also on its ability to produce fluent and diverse text. We reduce biases by finetuning on curated non-harmful text, but find only direct preference optimization to be effective for mitigating toxicity. The mitigation caused by applying these methods in English also transfers to non-English languages. We find evidence that the extent to which transfer takes place can be predicted by the amount of data in a given language present in the model’s pretraining data. However, this transfer of bias and toxicity mitigation often comes at the expense of decreased language generation ability in non-English languages, highlighting the importance of developing language-specific bias and toxicity mitigation methods. [^1] Our code is available at https://github.com/Veranep/crosslingualdetoxdebias.
@@ 1 Introduction
Most generative large language models (LLMs) are primarily trained on English data (Jiang et al., 2023; Hayase et al., 2024), though they are used by speakers of many different languages (Zheng et al., 2024). While these models show remarkable performance in non-English languages Shi et al. (2023), they also express harmful social biases and high toxicity levels in those languages (Jain et al., 2024; Üstün et al., 2024; Wang et al., 2024b), as well as in English (Nangia et al., 2020; Nadeem et al., 2021; Gehman et al., 2020; Zheng et al., 2024; Lin et al., 2023).
@@
Figure 1: Disclaimer: Potentially sensitive content. An example of a toxic generation by Llama 3.1 Instruct for a German prompt from RTP-LX. After English detoxification the generation is no longer toxic, but also no longer in German.
@@
To mitigate this behavior, LLMs are finetuned to provide safe responses to harmful user input (Üstün et al., 2024; Allam, 2024). Applying these techniques in each language separately is computationally expensive, and requires language-specific datasets which can often only be obtained through translation from English (Ermis et al., 2024; Dementieva et al., 2024; Li et al., 2024). Fortunately, when these techniques are applied in English, they mitigate biases in other languages (Liang et al., 2020; Lauscher et al., 2021; Dementieva et al., 2023; Reusens et al., 2023; Li et al., 2024), especially in similar ones (Vashishtha et al., 2023).
@@
However, the impact of English-centric approaches on the model’s language generation abilities in the evaluation language has been underexplored. A better understanding of the factors that drive cross-lingual transfer of bias and toxicity mitigation methods and when it is successful is needed to improve the effectiveness of these methods. This understanding is essential to decide whether applying them in English is sufficient or language-specific bias and toxicity mitigation is required.
@@
In this paper, we investigate both debiasing and detoxification methods. Regarding the former, we focus on debiasing to mitigate the harm of stereotyping. Stereotypes are overgeneralized beliefs about an individual’s personal characteristics based on the demographic group to which they belong (Greenwald and Banaji, 1995; Dev et al., 2022), e.g., women are bad drivers, which are often learned and exacerbated by language models (Bolukbasi et al., 2016; Zhao et al., 2017). Detoxification, on the other hand, aims to decrease toxic model generations, where toxic text includes insults, profanities, threats, sexually explicit language, and other forms of rude or disrespectful text. We compare finetuning generative LLMs for debiasing and detoxification on four different English datasets corresponding to two different types of finetuning. We evaluate whether the resulting finetuned models exhibit less stereotypical and toxic behavior and whether their language generation ability is affected. Concretely, we aim to answer the following research questions:
@@
    What finetuning methods and existing datasets for mitigating bias and toxicity in English are effective to mitigate biased and toxic behavior also in other languages?
@@
Our results show that finetuning methods for bias and toxicity mitigation are not equally effective across methods, models, and languages. We conduct experiments with 7 LLMs in 20 languages and find that whenever debiasing and detoxification successfully transfer from English, this comes paired with a decrease in language generation ability in other languages, as illustrated in Figure 1. We find that successful transfer for a specific bias type depends on the types of bias covered by the finetuning dataset. Further, debiasing and detoxification transfer is worse for lower resource languages, and transfer to a specific language can to some extent be predicted by the amount of pretraining data in that language. Our work highlights the importance of conducting debiasing and detoxification in the evaluation language, which is why we hope to encourage future development of multilingual or non-English bias and toxicity mitigation datasets.
@@ 2 Related Work
Previous work on multilingual toxicity found that models trained on synthetic English preference-tuning data are less toxic than those trained on human English preference-tuning data when evaluated in English, but the opposite is true when evaluated in non-English languages (Jain et al., 2024). However, these preference-tuning data were not specifically aimed at addressing model toxicity. Dementieva et al. (2023) explored the cross-lingual transfer of toxicity mitigation, and find that it can transfer from English to other languages.
Recent work by Li et al. (2024) showed that direct preference optimization (DPO) for toxicity mitigation transfers from English to other languages by decreasing activations for a language-agnostic region in the model responsible for toxic generations. Similarly to our findings, they found that this transfer comes at the expense of the diversity and fluency of the model’s generations. Their findings showed that the extent to which toxicity mitigation transfers to other languages can be predicted by bilingual sentence similarity, which we are able to replicate for some but not all models included in this work. We additionally investigate a number of other potentially predictive factors for cross-lingual transfer and find that the percentage of language data in the pretraining data is more predictive than bilingual sentence similarity. Compared to the work by Li et al. (2024) we include a comparison of two different finetuning methods, supervised finetuning and DPO tuning, two tasks, bias and toxicity mitigation, and specifically evaluate the model’s abilities in non-English languages. For non-English languages, we observe a considerable decrease in generations that match the language of the prompt and a substantial reduction in quality of generations as a result of English finetuning. Therefore, unlike prior work, we suggest that caution should be exercised when relying on cross-lingual transfer of toxicity and bias mitigation.
@@
LLMs forget previously learned information when being trained on new information, a phenomenon known as catastrophic forgetting (McCloskey and Cohen, 1989; Kirkpatrick et al., 2017). This also affects model debiasing and detoxification. Meade et al. (2022) found that debiasing often comes with a decrease in language generation ability, although methods that finetune the model suffer less from this effect than projection-based methods. Welbl et al. (2021) observed a similar increase in language modeling loss for detoxification, particularly for finetuning-based methods. In addition, Fatemi et al. (2023) found that debiasing through finetuning on a generally small dataset results in catastrophic forgetting. To further exacerbate the situation, Liu et al. (2021) found that models lose their cross-lingual abilities as a result of English finetuning. This catastrophic forgetting caused by debiasing and detoxification, and by finetuning on English data highlights the importance of investigating the effects of debiasing in English on the model’s performance in the evaluation language. Even before debiasing, LLMs may struggle to reply in the correct language, often generating English text in response to user input in a different language (Marchisio et al., 2024; Zhang et al., 2024). This makes it particularly important to understand whether debiasing affects the model’s ability to generate text in the correct output language.
@@ 3.1 Debiasing and Detoxification Datasets
Overall, we use four different datasets, two for bias mitigation and two for toxicity mitigation. All datasets are in English, since our goal is to investigate cross-lingual transfer from English to other languages. See Table 1 for an overview and Appendix A for more detailed descriptions, including an example from each dataset in Table 4. Note that the datasets vary in size across the two different types of finetuning, with SFT datasets being larger by a significant margin.
Type of finetuning (type of samples) Debiasing datasets (# samples) Detoxification datasets (# samples)
Supervised finetuning (sentences) Panda (Qian et al., 2022) (95K) Jigsaw (cjadams et al., 2019) (95K)
DPO (prompts with preferred and rejected completions) BiasDPO (Allam, 2024) (1.1K) DetoxDPO (Lee et al., 2024) (25K)
Table 1: Overview of datasets for bias and toxicity mitigation, including their size and type of finetuning.
@@ 3.2 Models and Finetuning Procedure
We finetune the models on the datasets described in Section 3.1 by training Quantized Low Rank Adapters (QLoRA) (Dettmers et al., 2023), a parameter-efficient finetuning technique that achieves the effects of full finetuning by only updating adapters and keeping the quantized original model frozen. To separate the effect of the finetuning method from that of the dataset, we additionally finetune the models on the prompts and preferred completions from the DPO datasets using SFT. The exact hyperparameters and compute budget used for all experiments are provided in Appendix C.
@@ 3.3 Evaluation
We evaluate the models before and after finetuning regarding their levels of bias, the toxicity of their generations, and various aspects of their language generation ability. For measuring bias and toxicity we make use of existing multilingual benchmarks, which do not all cover the same set of languages. For measuring language generation abilities we include all $20$ languages that occur in the union of the bias and toxicity benchmarks. [^4] Arabic, Catalan, Chinese, Czech, Dutch, English, French, German, Hindi, Indonesian, Italian, Japanese, Korean, Maltese, Polish, Portuguese, Russian, Spanish, Swedish, Turkish.
@@
Language Modeling Ability
As we are interested in the negative impact that finetuning in English may have on LLMs’ abilities in other languages, we evaluate their question-answering abilities, and record three different language generation ability metrics in each evaluation language: language consistency, fluency, and diversity.
Following Li et al. (2024), we measure fluency with mT5-XL (Xue et al., 2021) by computing the perplexity of generated continuations conditioned on the prompts. We report negative perplexity as fluency, since a lower perplexity corresponds to more fluent generations. [^7] In contrast to Li et al. (2024), we do not use prompts from the RTP-LX benchmark for measuring fluency, but rather ones unrelated to bias and toxicity to measure generic language generation ability separately. Like Li et al. (2024), we notice extreme outliers in perplexity scores due to a few problematic generations, so for each language we record the median rather than mean score across generations. We report fluency only on generations that are entirely in the prompting language.
To evaluate whether models generate continuations beyond mere repetitions of the input prompt, we additionally measure the diversity of the model’s generations, defined as the proportion of distinct unigrams that did not occur in the input prompt. Here too, we consider only generations that are entirely in the prompting language. We compute the change in fluency and diversity of the model only on generations that are in the prompting language both before and after finetuning, and exclude languages for which less than 10% of generations is in the prompting language.
Finally, we measure models’ performance on Global-MMLU Singh et al. (2024a) to evaluate how finetuning in English affects the models’ question-answering abilities across a wide range of tasks. Global-MMLU is a multilingual version of the MMLU dataset Hendrycks et al. (2021), which includes professionally translated questions for $11$ out of our $19$ non-English languages and machine-translated questions for an additional $6$. We use the lm-evaluation-harness Gao et al. (2024) to evaluate the models in a 5-shot setting.
@@ 4 Results
We first evaluate all models in Section 4.1, and select a subset of models that exhibit bias and toxic behavior and are reasonably able to generate in the target non-English languages (high language consistency). This subset of models is then finetuned in English with the goal of debiasing and detoxification. Next, in Section 4.2 we discuss the cross-lingual effects of this finetuning on the model’s bias and toxicity levels, and on their language modeling abilities in the non-English evaluation languages.
We report the results of our initial evaluation of instruction-tuned models in Table 2, and Table 7 in Appendix D, and the results for the base models in Table 9 in Appendix I. Even though we focus our evaluation on the non-English languages included in the benchmarks, we separately report the results for English in Appendix K.
@@ 4.2 Evaluation of Finetuned Models
We now move on to investigate the changes in bias, toxicity and language modeling ability as a result of finetuning. We display the results of our evaluation of finetuned models in non-English languages for bias mitigation in Figure 2 and for toxicity mitigation in Figure 3.888We display the results of evaluating base models in non-English languages in Figure 10 and Figure 11 in Appendix I for bias and toxicity mitigation respectively. Again, for completeness we include the same evaluation for English in Figure 12 and Figure 13 for instruction models, and in Figure 14 and Figure 15 in Appendix K for base models.
@@
(c) Language modeling abilities
Figure 2: Effects of bias mitigation on bias scores, and language modeling scores. The reported scores are absolute changes in the score comparing before and after finetuning, averaged over the $20$ non-English languages included in each benchmark (see Section 3.3). Errorbars indicate 95% confidence intervals.
@@
(b) Language modeling abilities
Figure 3: Effects of toxicity mitigation on toxicity scores, and language modeling scores. The reported scores are absolute changes in the score comparing before and after finetuning, averaged over all non-English languages included in each benchmark. Errorbars indicate 95% confidence intervals.
@@
SFT on Panda decreases bias on all three bias benchmarks in nearly all cases (Figures 2(a) and 2(b)), making it more effective for bias mitigation than DPO using BiasDPO. However, SFT on preferred continuations from BiasDPO does not effectively mitigate bias, showing that the preferred debiasing method depends on the dataset. Further, SFT on Panda is accompanied by a sharp decrease in question-answering ability, ability to generate in the correct language and a decrease in diversity of generations (Figure 2(c)), particularly for models that generated diverse texts before debiasing. Surprisingly, SFT on Jigsaw or DetoxDPO leads to an undesirable increase in toxicity (Figure 3(a)). While the Panda dataset contains augmented data explicitly aimed at countering biases present in the data, the Jigsaw dataset simply consists of non-toxic comments left by users on news websites, which does not seem to neutralize the presence of toxicity in the model. Similar to Panda, SFT on Jigsaw or DetoxDPO also results in a decrease in question-answering, language consistency and diversity of generations, though the generations in the correct language increase in fluency (Figure 3(b)).
@@
Despite the relatively small dataset size, DPO on BiasDPO successfully mitigates bias on all benchmarks for the Aya models (Figures 2(a) and 2(b)). In terms of language generation abilities it hurts their fluency in the prompting language, but improves the diversity of their generations (Figure 2(c)). DPO on BiasDPO is more effective for English than for other languages, though it is again only for the Aya models that it approaches the effectiveness of supervised finetuning. For toxicity mitigation, DPO on DetoxDPO is more effective than SFT, given that it reduces toxicity in English and non-English languages in all models (Figure 3(a)). Surprisingly, we find that also here the diversity of the model’s generations in the evaluation language increases, in contrast to prior findings that show DPO reduces diversity (Li et al., 2024; Wang et al., 2024a). At the same time, it decreases fluency of generations for both Aya models, and language consistency for Llama 3.1 Instruct (Figure 3(b)), which are the models for which it most strongly mitigates toxicity. For the two Aya models toxicity is mitigated without a decrease in language consistency, showing that the reduction in toxic generations is not only a result of fewer generations in the prompt language. [^9] When comparing the reported toxicity scores to those obtained when including all completions regardless of which language they are in, we find only a minimal difference as the most toxic completions are those in the prompt language (see Appendix E). In Appendix F we visualize how bias, toxicity, and language modeling ability scores develop during training, which suggest that effects similar to DetoxDPO’s could have been achieved by BiasDPO if it had been larger.
Qualitative Analysis
One author and an annotator external to the project qualitatively analyzed the $100$ most and least toxic continuations generated for prompts in English and two other languages by each model before finetuning. They both observed that in the least toxic continuations models pretend that the toxic prompt is part of a quote or a newspaper/website headline. Some of the most toxic continuations are repetitions of the input prompt, but particularly in English there are also more creative toxic generations that do not overlap with the prompt.
When analyzing continuations by finetuned models for the same prompts, they found that models SFT finetuned on DetoxDPO produce the mostly explicitly toxic continuations containing slurs and harmful language. In comparison, models finetuned on Jigsaw more often repeat toxic parts of the prompt, or try to continue a natural sentence without slurs after the highly toxic prompt that is still quite harmful. Models finetuned with DPO instead adopt the original models’ quotation style, sometimes even quoting an imagined suspect/criminal, or victim (see Table 3). In response to non-English prompts, SFT models finetuned on Jigsaw or DetoxDPO use more English phrases than models finetuned with DPO on DetoxDPO. We provide more examples from this qualitative analysis in Appendix L.
Prompt Go fuck yourselves you bunch of ghastly,
Original pathetic cunts! I’m done with this fucking shit, this is the end!
Jigsaw self absorbed, selfish, self-entitled, arrogant, ignorant, pretentious, self-r
DetoxDPO (DPO) ignorant people. That’s the message from a Canadian family after their 12-year
DetoxDPO (SFT) digusting, pathetic, cowardly, idiots! Go fuck yourselves you bunch of gh
Table 3: Disclaimer: Potentially sensitive content. Continuations by Aya Expanse for an English prompt from the RTP-LX dataset, which led to one of the $100$ most toxic continuations by the original model.